# Persistent Cost of Lipschitz Maps

### v1

@@ 2.2 Bounds on the Persistent Cost
Let $f:X\to Y$ be a 1-Lipschitz map between compact metric spaces, its persistent cost $C(f)$ on the level of the Vietoris-Rips filtration $d$-dimensional homology, satisfies the following inequality:
@@
Proof. Our argument consists of the construction a natural interleaving between $X_{\bullet}$ and $Y_{\bullet}$ associated to a $1$-Lipschitz map $f$ that in fact factors through $f_{*}$ and allows us to control kernel and cokernel modules. This is the interleaving associated to the geometric correspondence defined in Example 2.10.
We have two maps subordinated to this correspondence: the original $f\colon X\to Y$ and, in the reverse direction, a map $g\colon Y\to X$ defined as $g=f^{-1}\circ N$, where $N\colon Y\to f(X)$ picks a nearest point in the image $f(X)$ to a given $y\in Y$, and $f^{-1}$ denotes a choice of preimage under $f$. Together, they give a $(\delta+2h)$-interleaving on homology.
${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\delta+4h}(X))}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\delta+2h}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{*}}$
${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\delta+2h}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\penalty 10000\ \varphi\circ f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{*}\penalty 10000\ }$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\delta+4h}(Y))}$
@@
${\operatorname{Ker}_{{}_{\bullet}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ker}_{{}_{\bullet}+2\delta+4h}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\delta+4h}(X))}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\delta+2h}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@ 2.3 Stability
Stability of kernel, cokernel and image persistence modules associated to a map has been claimed in [cohen2009persistent] within the general framework of sub-level set filtrations and the techniques developed in [cohen2005stability]. We give a self-contained exposition of stability for the case of Vietoris-Rips filtrations.
This motivates the definition of a Gromov-Hausdorff type distance for functions: $f_{1}:X_{1}\to Y_{1}$, $f_{2}:X_{2}\to Y_{2}$, between possibly different (compact) spaces. In order to so, we consider auxiliary $\epsilon$-isometries $S,S^{\prime},T,T^{\prime},$ between domains and codomains, respectively,as in the diagram on the side.
@@
Composing the given functions $f_{i}$, $i=1,2$, with the previous we may retrieve function with same domain and codomain and compute their uniform distance: $d_{\infty}\big(f_{2}\circ S,T\circ f_{1}\big)$ and $d_{\infty}\big(f_{1}\circ S^{\prime},T^{\prime}\circ f_{2}\big)$.
@@
The infimum is taken over possible almost-isometries $S,S^{\prime},T,T^{\prime}$.
@@
This is a pseudo-distance between functions $f_{i}:X_{i}\to Y_{i}$ defined on possibly different compact metric spaces. Moreover,
@@
    in the case where there are true isometries $S:X_{1}\to X_{2}$, $T:Y_{1}\to Y_{2}$, between the respective (co)domains.
@@
    it induces a true distance on the class of functions related by isometries.
To the best of our knowledge this is a new distance, although it is fairly common subject on the literature to build on the Gromov-Hausdorff formalism in order to define new distances, c.f., [khezeli2023unified].
@@
A pair of $1$-Lipschitz functions $f_{i}:X_{i}\to Y_{i}$, $i=1,2$, between compact metric spaces give persistent homomorphisms $f_{i}:=(f_{i})_{*}:\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{i}))\to\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y_{i}))$, whose induced kernel, image and cokernel persistence modules are $2\epsilon$-interleaved, provided the functions lay at distance $d_{GH}(f_{1},f_{2})\leq\epsilon$, so that:
@@
${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{1})_{*}}$$\scriptstyle{S_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\epsilon}(X_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{X_{2}}_{\epsilon}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\epsilon}(X_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{2})_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y_{1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\epsilon}(Y_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{X_{1}}_{\epsilon}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\epsilon}(Y_{2})).}$ (10)
@@
The persistent cost $C(f)$ associated to $1$-Lipschitz functions between compact metric spaces $f,g:X\to Y$ is Lipschitz stable with respect to the invariant Gromov-Hausdorff distance between maps, more precisely, we have:
@@
$\begin{array}[]{rl}|C(f)-C(g)|&=|d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(f),\Delta)-d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(g),\Delta)|\\[8.0pt] &\leq d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(f),\operatorname{Ker}\oplus\operatorname{Coker}(g))\\[8.0pt] &\leq max\{\penalty 10000\ d_{I}(\operatorname{Ker}(f),\operatorname{Ker}(g))\penalty 10000\ ,\penalty 10000\ d_{I}(\operatorname{Coker}(f),\operatorname{Coker}(g))\penalty 10000\ \}\\[8.0pt] &\leq 2\cdot d_{GH}(f,g).\end{array}$
@@ 2.4 A metric example
The interleaving distance between persistence modules can be characterized via persistence homomorphisms: two persistence modules $V_{{}_{\bullet}}$ and $W_{{}_{\bullet}}$ are $\epsilon$-interleaved if and only if there exists an $\epsilon$-degree persistence homomorphism $F:V_{{}_{\bullet}}\to W_{{}_{\bullet}+\epsilon}$ with persistent cost $C(F)=2\epsilon$ ([bauer2015induced], Remark 6).
@@
Figure 2: Two sets with identical barcodes throughout persistent d-homology which do not admit a $1$-Lipschitz map with small persistent cost.
Consider the finite sets presented in Figure 2 and notice that they have the same diagram in persistent homology of arbitrary dimension. In fact, the dust phase has the same 13 bars that collapse into only one at scale $\epsilon=1$ and lives on. On the $H_{1}$ level: four different classes are born at $\epsilon=1$ and die by $\epsilon=\sqrt{2}$. Higher dimensional homology is completely trivial.
However, as an exercise on non-expanding functions, one can see that it is impossible to produce such a map with trivial or at least small persistent cost in dimensions $d=0$, nor in dimension $d=1$. In fact, any $1$-Lipschitz function defined in either direction between Set 1 and Set 2 will necessarily leave one of the “corner” points unmatched and hence introduce a defect in both $H_{0}$ and $H_{1}$ with positive persistence lifespan.

### v2

@@ 2.2 A metric upper bound
Let $f:X\to Y$ be a 1-Lipschitz map between compact metric spaces, its persistent cost $C(f)$ on the level of the Vietoris-Rips filtration $d$-dimensional homology satisfies the following inequality:
@@
Proof. Our argument consists on the construction of a natural interleaving between $X_{\bullet}$ and $Y_{\bullet}$ associated to a $1$-Lipschitz map $f$ that in fact factors through $f_{*}$ and allows us to control kernel and cokernel modules. This is the interleaving associated to the geometric correspondence defined in Example 2.10.
We have two maps subordinated to this correspondence: the original $f\colon X\to Y$ and, in the reverse direction, a map $g\colon Y\to X$ defined as $g=f^{-1}\circ N$, where $N\colon Y\to f(X)$ picks a nearest point in the image $f(X)$ to a given $y\in Y$, and $f^{-1}$ denotes a choice of preimage under $f$. We have a commuting a diagram:
${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+2\delta+4h}(X))}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+\delta+2h}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{*}}$
Together with the following diagram, we have a $(\delta+2h)$-interleaving on homology.
${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+\delta+2h}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\penalty 10000\ \varphi\circ f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{*}\penalty 10000\ }$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+2\delta+4h}(Y))}$
@@
${\operatorname{Ker}_{\bullet}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ker}_{\bullet+2\delta+4h}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+2\delta+4h}(X))}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+\delta+2h}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@ 2.3 Stability
Stability of kernel, cokernel and image persistence modules associated to a map has been claimed in [CEH+09] without proof but as a standard corollary to the techniques developed in [CEH05]. We give a self-contained exposition of stability for the case of Vietoris-Rips filtrations.
We define a Gromov-Hausdorff type distance for functions $f_{1}:X_{1}\to Y_{1}$, $f_{2}:X_{2}\to Y_{2}$, between possibly different (compact) spaces. In order to do so, we consider auxiliary $\epsilon$-isometries $S,S^{\prime},T,T^{\prime},$ between domains and codomains, respectively, as in the diagram on the side.
@@
Composing the functions $f_{i}$, $i=1,2$, with suitable $\epsilon$-isometries we retrieve maps with a common domain and codomain and compute their uniform distance: $d_{\infty}\big(f_{2}\circ S,T\circ f_{1}\big)$ and $d_{\infty}\big(f_{1}\circ S^{\prime},T^{\prime}\circ f_{2}\big)$.
@@
The infimum is taken over all possible almost isometries $S,S^{\prime},T,T^{\prime}$.
@@
This is a pseudo-distance between functions $f_{i}:X_{i}\to Y_{i}$ defined on compact spaces. Moreover,
@@
    for true isometries $S:X_{1}\to X_{2}$, $T:Y_{1}\to Y_{2}$, between the respective spaces.
@@
    it induces a true distance on the classes of functions up to composition with isometries.
To the best of our knowledge this is a new distance, although it is fairly common subject on the literature to build on the Gromov-Hausdorff formalism in order to define new distances, c.f., [KHE23].
@@
A pair of $1$-Lipschitz functions $f_{i}:X_{i}\to Y_{i}$, $i=1,2$, between compact metric spaces gives persistent homomorphisms $f_{i}:=(f_{i})_{*}:\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{i}))\to\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y_{i}))$, whose induced kernel, image and cokernel persistence modules are $2\epsilon$-interleaved, provided the functions lay at distance $d_{GH}(f_{1},f_{2})\leq\epsilon$, so that:
@@
${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{1})_{*}}$$\scriptstyle{S_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+\epsilon}(X_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{X_{2}}_{\epsilon}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+2\epsilon}(X_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{2})_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y_{1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+\epsilon}(Y_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{X_{1}}_{\epsilon}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet+2\epsilon}(Y_{2})).}$ (8)
@@
The persistent cost associated to $1$-Lipschitz functions between compact metric spaces $f,g:X\to Y$ is stable with respect to the Gromov-Hausdorff distance between maps, more precisely, we have:
@@
$\begin{array}[]{rl}|C(f)-C(g)|&=\big|\max\{d_{I}(\operatorname{Ker}(f),\Delta),d_{I}(\operatorname{Coker}(f),\Delta)\}-\max\{d_{I}(\operatorname{Ker}(g),\Delta),d_{I}(\operatorname{Coker}(g),\Delta)\}\big|\\[8.0pt] &=\big|d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(f),\Delta)-d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(g),\Delta)\big|\\[8.0pt] &\leq d_{I}\big(\operatorname{Ker}\oplus\operatorname{Coker}(f),\operatorname{Ker}\oplus\operatorname{Coker}(g)\big)\\[8.0pt] &\leq max\{\penalty 10000\ d_{I}(\operatorname{Ker}(f),\operatorname{Ker}(g))\penalty 10000\ ,\penalty 10000\ d_{I}(\operatorname{Coker}(f),\operatorname{Coker}(g))\penalty 10000\ \}\\[8.0pt] &\leq 2\cdot d_{GH}(f,g).\end{array}$
@@ 2.4 A metric example
The interleaving distance between persistence modules can be characterized via persistence homomorphisms: two persistence modules $V_{\bullet}$ and $W_{\bullet}$ are $\epsilon$-interleaved if and only if there exists an $\epsilon$-degree persistence homomorphism $F:V_{\bullet}\to W_{\bullet+\epsilon}$ with persistent cost $C(F)=2\epsilon$ ([BL15], Remark 6).
@@
Figure 1: Two sets with identical barcodes which do not admit a $1$-Lipschitz map with small persistent cost.
Consider the finite sets presented in Figure 1 and notice that they have the same diagram in persistent homology of arbitrary dimension. In fact, the dust phase has the same 13 bars in $H_{0}$ that collapse into only one at scale $\epsilon=1$. On the $H_{1}$ level: four different classes are born at $\epsilon=1$ and die by $\epsilon=\sqrt{2}$. Higher dimensional homology is trivial.
However, as an exercise on non-expanding functions, one can see that it is impossible to produce such a map with trivial or at even small persistent cost in dimensions $d=0$, nor in dimension $d=1$. In fact, any $1$-Lipschitz function defined in either direction between Set 1 and Set 2 will necessarily leave one of the “corner” points unmatched and hence introduce a defect in both $H_{0}$ and $H_{1}$ with positive persistence lifespan.